OpenStudy (anonymous):

Use the quadratic formula to solve the equation below. -6x= -4x^2 +18 A. X=3 | X= -1.5 B.X=1.5 | X= -3 C.X=3 | X= -3 D.X=-3 | X= -1.5

OpenStudy (anonymous):

Do you think you might be able to help? @ash2326

OpenStudy (ash2326):

Standard form of a quadratic equation \[ax^2+bx+c=0\] here we have \[-6x=-4x^2+18\] or \[4x^2-6x-18=0\] Do you get this part? I've just rearranged the equation :)

OpenStudy (anonymous):

Yes.

OpenStudy (ash2326):

The solution of quadratic equation is \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] here a=4 b=-6 and c=-18 Can you plug in the values and solve for x?

OpenStudy (anonymous):

Can you help me on how to find x. I don't quite get it.

OpenStudy (ash2326):

ok :)

OpenStudy (ash2326):

\[x=\frac{-(-6)\pm \sqrt{(-6)^2-4\times 4 \times (-18)}}{2\times 4}\] \[x=\frac{ -(-6)\pm \sqrt{36+288}}{2\times 4}\] \[x=\frac{ -(-6)\pm \sqrt{324}}{2\times 4}\] \[x=\frac{ 6\pm \sqrt{324}}{2\times 4}\] \[x=\frac{6\pm 18}{8}\] so \[x=\frac{6+18}{8}, \frac{6-18}{8} \] Just simplify now :)

OpenStudy (anonymous):

So i just did 6+18=24 and then divided it by 8 and got 3

OpenStudy (ash2326):

yes, find the other one also

OpenStudy (anonymous):

Then I did 6-18= -12 and -12 divided by 8 and got -1.5

OpenStudy (anonymous):

So it would be A

OpenStudy (ash2326):

correct :) Did you understand?

OpenStudy (anonymous):

Yeah. Better then I did before

OpenStudy (anonymous):

Thank you :) I have a few more if you wouldn't mind helping me

OpenStudy (ash2326):

I'm here for some time, I can help :P|)